OpenStudy (anonymous):

Using the law of cosines, Find c. Triangle ABC, a=8, b=6, y=60

OpenStudy (anonymous):

what is y?

OpenStudy (anonymous):

an angle?

OpenStudy (anonymous):

is angle y opposite side c?

OpenStudy (anonymous):

\(c=a^2+b^2-2ab\cos(C)\) is how this is usually written. my guess is that you have \[c^2=8^2+6^2-2\times 8\times 6\cos(60)\]

OpenStudy (anonymous):

the triangle actually does not show where y is

OpenStudy (anonymous):

since \(\cos(60)=\frac{1}{2}\) you get \[c^2=8^2+8^2-8\times 6\] to calculate

OpenStudy (anonymous):

sorry i meant \[c^2=8^2+6^2-6\times 8\]

OpenStudy (anonymous):

80?

OpenStudy (anonymous):

if you are supposed to use the law of cosines, this is the only thing it can be

OpenStudy (anonymous):

i get \(c^2=52\) making \(c=\sqrt{52}\)

OpenStudy (anonymous):

52

OpenStudy (anonymous):

ok, thanks!

OpenStudy (anonymous):

or \(c=2\sqrt{13}\) if you prefer